OpenStudy (anonymous):

Explain why the degree of the term 4^5 is not 5. What is the degree?

OpenStudy (mertsj):

Because the degree of a term is the exponent of the variable.

OpenStudy (anonymous):

okay thanks and what would the degree be? @Mertsj

OpenStudy (mertsj):

Does your term have a variable?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

What do you think the exponent of the variable was that made it possible for the variable to be unwritten?

OpenStudy (anonymous):

uhh 0

OpenStudy (anonymous):

im confused

OpenStudy (mertsj):

And that is the degree of your term...all constants have degree 0

OpenStudy (anonymous):

oooh thanks!!

OpenStudy (mertsj):

\[4^5x^0=4^5(1)=4^5=1024\]

OpenStudy (mertsj):

yw